OpenStudy (anonymous):

write the equation of the line passing through points (6,-3), (-2,-3)

OpenStudy (anonymous):

Not completely sure if this is correct but the first thing I did was found the slope of the two points. The slope is 0. I then put it into y = mx + b form. So I have y = 0x + b. I used the first point to figure out the b. -3 = 0(6) + b. -3 = 0 + b. Subtract 0 from both sides, b = -3. So the final equation is y = 0x -3 *I think*